### v1

@@ 1 Introduction
To address these challenges, we propose a novel framework, Prompt Learning with Integrated Label-Free Test-Time Adaptation for ZerO-ShoT Anomaly Detection (PILOT), a framework designed to mitigate the limitations of current ZSAD methods under domain shift. Instead of relying on a single learnable prompt, PILOT constructs a pool of learnable prompts, allowing the prompts to adaptively learn to various anomalies present in the auxiliary dataset. This design enables the model to adaptively weight the most relevant prompts for each input image, allowing for more nuanced representations of diverse anomalies and supporting more effective TTA. During inference, our label-free test-time adaptation strategy updates all learnable prompts according to their alignment with each test image, using high-confidence pseudo-labels generated from unlabeled target data. This targeted adaptation facilitates rapid and reliable adjustment to new domains (i.e., the test domain) without requiring any labeled data. In addition, to mitigate overfitting to the auxiliary dataset and enhance adaptation stability, PILOT incorporates an attribute memory bank, which consists of a curated set of fixed prompts that describe normal and anomalous states, leveraging pretrained semantic knowledge from the text encoder. This attribute memory bank serves as a semantic anchor, helping to stabilize localization performance during TTA. Together, these innovations enable PILOT to achieve robust anomaly detection and localization across a variety of challenging real-world scenarios. Our contributions are summarized as follows:
@@ 2 Proposed Method
ZSAD involves detecting and localizing anomalies in images from categories not encountered during the training phase. Formally, in the training phase, we employ an auxiliary dataset defined as $\mathcal{D}_{\mathrm{aux}}={(I_{i},y_{i},G_{i})}_{i=1}^{|\mathcal{D}_{\mathrm{aux}}|}$, where each image $I_{i}\!\in\!\mathbb{R}^{H\times W\times 3}$ is paired with an image-level label $y_{i}\!\in\!\{0,1\}$ (0 denotes normal and 1 denotes anomalous) and a pixel-level anomaly mask $G_{i}\!\in\!\{0,1\}^{H\times W}$. In the inference phase, we consider an unlabeled target dataset $\mathcal{D}_{\mathrm{tar}}=\bigl{\{}I_{t}\bigr{\}}_{t=1}^{|\mathcal{D}_{\mathrm{tar}}|},$ where each $I_{t}\in\mathbb{R}^{H\times W\times 3}$. For every $I_{t}$, the model outputs an anomaly score $\hat{y}_{t}\in[0,1],$ (with values closer to 1 indicating higher abnormality likelihood), along with a $\hat{G}_{t}\in[0,1]^{H\times W}$ that spatially localizes potential anomalous regions at pixel level. Crucially, the category sets of $\mathcal{D}_{\mathrm{aux}}$ and $\mathcal{D}_{\mathrm{tar}}$ are strictly disjoint, ensuring that the target domain remains unseen during training and thus establishing a true zero-shot setting.
@@ 2.1 Learnable Prompt Pool
[…]
@@ 2.3 Prompt Aggregation via Orthogonal Projection
Following [Yang et al.(2021)Yang, He, Fan, Shi, Xue, Li, Ding, and Huang], to robustly leverage both the stable semantics from the attribute embedding $T_{att}$ and the adaptive semantics from the learnable prompts ($T_{n}$ for normal and $T_{a}$ for anomalous), we aggregate these representations using an orthogonal projection operation instead of simple averaging or concatenation. For each prompt embedding $T_{(\cdot)}$ (where ()$(\cdot)\in\{n,a\}$), we decompose it into components parallel and orthogonal to $T_{att}$:
@@ 3 Experiments
Table 2: Comparison of PILOT vs. state-of-the-art on industrial benchmarks. $\mathbf{Img}$ and $\mathbf{Pix}$ denote image-level and pixel-level metrics, respectively. “w/o TTA” and “w/ TTA” indicate performance without and with TTA, respectively. Arrows indicate change in performance: $\uparrow$(increase >0.5>0.5> 0.5), $\downarrow$(decrease <0.5<-0.5< - 0.5), and $\circ$(no significant change, ±0.5$\pm$ 0.5± 0.5).
Dataset APRIL-GAN AnomalyCLIP AdaCLIP PILOP w/o TTA w/ TTA w/o TTA w/ TTA w/o TTA w/ TTA w/o TTA w/ TTA MVTec AD $\mathbf{Img}$ (86.5, 93.2) (82.9$\downarrow$, 89.8$\downarrow$) (91.6, 96.3) (89.2$\downarrow$, 95.0$\downarrow$) (91.3, 95.8) (90.1$\downarrow$, 95.8$\circ$) (90.2, 95.7) (92.1$\uparrow$, 96.6$\uparrow$) $\mathbf{Pix}$ (87.6, 44.0) (86.0$\downarrow$, 42.0$\downarrow$) (90.7, 78.7) (88.1$\downarrow$, 73.5$\downarrow$) (89.4, 37.8) (87.0$\downarrow$, 29.5$\downarrow$) (89.4, 79.2) (90.3$\uparrow$, 80.2$\uparrow$) VisA $\mathbf{Img}$ (78.2, 81.3) (84.5$\uparrow$, 88.0$\uparrow$) (81.8, 85.1) (84.3$\uparrow$, 85.8$\uparrow$) (81.6, 84.1) (83.0$\uparrow$, 85.0$\uparrow$) (82.0, 84.1) (84.1$\uparrow$, 85.7$\uparrow$) $\mathbf{Pix}$ (94.2, 86.8) (92.0$\downarrow$, 84.0$\downarrow$) (95.3, 85.1) (85.5$\downarrow$, 82.6$\downarrow$) (95.5, 77.8) (93.8$\downarrow$, 72.6$\downarrow$) (94.8, 85.6) (96.0$\uparrow$, 87.2$\uparrow$) BTAD $\mathbf{Img}$ (73.8, 68.9) (76.9$\uparrow$, 73.2$\uparrow$) (83.8, 88.4) (85.7$\uparrow$, 93.0$\uparrow$) (90.4, 90.7) (91.6$\uparrow$, 92.5$\uparrow$) (90.0, 91.6) (95.3$\uparrow$, 96.8$\uparrow$) $\mathbf{Pix}$ (60.8, 25.0) (58.0$\downarrow$, 23.0$\downarrow$) (94.4, 73.6) (94.1$\circ$, 72.0$\downarrow$) (94.8, 32.5) (93.0$\downarrow$, 25.5$\downarrow$) (94.9, 73.0) (96.3$\uparrow$, 74.9$\uparrow$) SDD $\mathbf{Img}$ (80.0, 71.7) (79.9$\circ$, 71.6$\circ$) (85.5, 81.7) (85.7$\circ$, 81.8$\circ$) (81.4, 72.8) (83.0$\uparrow$, 74.0$\uparrow$) (87.3, 85.2) (87.9$\uparrow$, 85.7$\uparrow$) $\mathbf{Pix}$ (79.8, 65.1) (77.5$\downarrow$, 63.0$\downarrow$) (90.7, 66.6) (89.9$\downarrow$, 63.8$\downarrow$) (71.7, 17.6) (66.7$\downarrow$, 12.5$\downarrow$) (88.7, 69.6) (92.4$\uparrow$, 73.4$\uparrow$) DAGM $\mathbf{Img}$ (94.6, 84.0) (95.0$\circ$, 84.5$\circ$) (96.1, 90.2) (97.3$\uparrow$, 92.0$\uparrow$) (97.3, 93.1) (96.8$\circ$, 92.2$\circ$) (97.1, 91.8) (98.5$\uparrow$, 92.8$\uparrow$) $\mathbf{Pix}$ (82.4, 66.2) (80.5$\downarrow$, 64.0$\downarrow$) (95.6, 91.0) (93.1$\downarrow$, 90.2$\downarrow$) (91.1, 62.3) (90.2$\downarrow$, 58.8$\downarrow$) (95.8, 91.8) (96.0$\circ$, 91.4$\circ$) DTD-Synthetic $\mathbf{Img}$ (86.7, 95.2) (85.0$\downarrow$, 93.5$\downarrow$) (93.7, 97.3) (93.2$\circ$, 96.9$\circ$) (97.5, 98.2) (97.5$\circ$, 98.4$\circ$) (95.3, 98.0) (97.7$\uparrow$, 98.4$\uparrow$) $\mathbf{Pix}$ (95.3, 86.9) (93.5$\downarrow$, 84.5$\downarrow$) (97.8, 91.1) (93.3$\downarrow$, 83.4$\downarrow$) (98.2, 86.5) (94.9$\downarrow$, 73.3$\downarrow$) (97.3, 89.9) (98.0$\uparrow$, 90.2$\circ$)
@@ 3.1 Comparison with State-of-the-Art
As reported in Table 5, PILOT establishes state-of-the-art performance across all benchmarks for both image-level anomaly detection and pixel-level localization in the zero-shot setting. This can be benefited from PILOP comprehensive approach in addressing critical challenges prevalent in ZSAD, that explicitly tackles domain shift in test phase. Prior approaches [Cao et al.(2023b)Cao, Xu, Sun, Cheng, Du, Gao, and Shen, Jeong et al.(2023)Jeong, Zou, Kim, Zhang, Ravichandran, and Dabeer] rely solely on predefined prompts that may not generalize well to unseen domains, while method such as [Chen et al.(2023)Chen, Han, and Zhang, Cao et al.(2024)Cao, Zhang, Frittoli, Cheng, Shen, and Boracchi, Zhou et al.(2022a)Zhou, Yang, Loy, and Liu, Zhou et al.(2024)Zhou, Pang, Tian, He, and Chen] utilized lightweight visual adapters or single learnable prompt to capture anomaly cues, but tend to overfit to the auxiliary training set, limiting their robustness under distribution shift. PILOT distinguishes itself by dynamically identifying the most relevant prompts for each input and refining only those learnable prompt parameters during inference using high-confidence pseudo-labels from the unlabeled target data. Note that while existing baselines can utilize the proposed image-level TTA strategies, these often degrade pixel-level localization performance and do not yield consistent improvements across both detection and localization tasks. On the contrary, PILOT uniquely achieves gains in both image-level and pixel-level metrics under TTA, a phenomenon analyzed in detail in Section 3.2. We also compare against state-of-the-art full-shot methods in Appendix F.1, which show that PILOT achieves image-level accuracy on par.
@@ 3.3 Module Ablation
SLP (90.8, 91.6) (91.4$\uparrow$, 91.9$\circ$) (94.2, 82.1) (90.7$\downarrow$, 79.8$\downarrow$)
+Attr (90.5, 91.3) (91.2$\uparrow$, 91.7$\circ$) (93.9, 81.5) (94.0$\circ$, 81.5$\circ$)
Full (90.3, 91.1) (92.6$\uparrow$, 92.7$\uparrow$) (93.5, 81.5) (94.8$\uparrow$, 82.9$\uparrow$)
@@ 4 Conclusion
This appendix provides supplementary details for the BMVC 2025 paper titled ”Zero-Shot Anomaly Detection with Dual-Branch Prompt Learning”.
@@ B.2 Domain Discrepancy Analysis
To validate that MVTec AD and VisA indeed represent statistically distinct domains suitable for strict zero-shot evaluation, we quantitatively assessed the distributional discrepancy between their CLIP global feature embeddings. Specifically, we employed Maximum Mean Discrepancy (MMD2[Gretton et al.(2012)Gretton, Borgwardt, Rasch, Schölkopf, and Smola], Fréchet Inception Distance (FID) [Heusel et al.(2017)Heusel, Ramsauer, Unterthiner, Nessler, and Hochreiter], and the Kolmogorov–Smirnov (KS) [Massey Jr(1951)] test. Our analysis yielded $\mathrm{MMD}^{2}=0.0328$ ($p$-value =0.0099=0.0099= 0.0099), indicating a significant difference in embedding distributions. The FID score between the two datasets is 51.951.951.9, further corroborating a substantial distributional gap. Additionally, the KS test computed across all feature dimensions revealed that the minimum, mean, and maximum $p$-values are 0, $5.16\times 10^{-5}$, and 0.03290.03290.0329, respectively, suggesting statistically significant differences in nearly every dimension of the embedding space. These results collectively confirm that MVTec AD and VisA are non-overlapping in the learned representation space, thus justifying their use as auxiliary and target datasets in our zero-shot anomaly detection experiments.
@@ D.1 PILOT Details
[…]
@@ E.2 Pseudo label Statistic
[…]
@@
E.3 Ablation Study on ohter TTA Method
@@ E.7 Prompt Aggregation Ablation
Proj. Extra Image Pixel w/o TTA w/ TTA w/o TTA w/ TTA GF $\checkmark$ (91.2, 90.0) (91.5$\circ$, 90.8$\uparrow$) (88.6, 77.8) (89.4$\uparrow$, 77.6$\circ$) CA $\checkmark$ (92.1, 92.4) (93.1$\uparrow$, 92.9$\uparrow$) (93.7, 81.3) (86.4$\downarrow$, 75.3$\downarrow$) OP $\times$ (90.3, 91.1) (92.6$\uparrow$, 92.7$\uparrow$) (93.5, 81.5) (94.8$\uparrow$, 82.9$\uparrow$)
@@ Appendix G Medical anomaly localization results
Medical anomaly localization results are reported in Table 14, with all evaluations conducted without TTA. Since medical datasets lack explicit “normal” examples, our TTA strategy cannot be applied; nevertheless, PLIOT still produces competitive localization maps, achieving the highest mean AUROC and the second-best AUPRO across benchmarks. These results underscore the generality of our dual-branch prompt learning framework: even in the absence of adaptation, dynamically weighting prompts with the most relevant anomaly cues deliver state-of-the-art performance in medical domains.
@@ Appendix I Compared with AnovL
Table 16: Comparison between PILOT and AnovL on two benchmarks (MVTec AD and VisA). “w/o TTA” and “w/ TTA” indicate performance without and with test-time adaptation, respectively. Arrows denote change under TTA: $\uparrow$(improvement), $\downarrow$(decline), $\circ$(no significant change, ±0.5$\pm$ 0.5± 0.5).
Method MVTec AD VisA Image (AUROC, AP) Pixel (AUROC, AUPRO) Image (AUROC, AP) Pixel (AUROC, AUPRO) w/o TTA w/ TTA w/o TTA w/ TTA w/o TTA w/ TTA w/o TTA w/ TTA AnoVL (87.9, 93.8) (87.9$\circ$, 93.8$\circ$) (82.7, 60.2) (84.8$\uparrow$, 68.4$\uparrow$) (72.6, 76.9) (72.6$\circ$, 76.9$\circ$) (84.4, 62.0) (85.6$\uparrow$, 64.1$\uparrow$) PILOT (90.2, 95.7) (92.1$\uparrow$, 96.6$\uparrow$) (89.4, 79.2) (90.3$\uparrow$, 80.2$\uparrow$) (82.0, 84.1) (84.1$\uparrow$, 85.7$\uparrow$) (94.8, 85.6) (96.0$\uparrow$, 87.2$\uparrow$) + AnoVL Loss (90.2, 95.7) (87.5$\downarrow$, 89.7$\downarrow$) (89.4, 79.2) (66.4$\downarrow$, 52.3$\downarrow$) (82.0, 84.1) (73.9$\downarrow$, 78.1$\downarrow$) (94.8, 85.6) (79.4$\downarrow$, 65.8$\downarrow$)
@@
4:   $x_{\rm cls},\{X^{(m)}\}_{m\in M}\leftarrow f_{visual}(I)$triangleright extract global & patch features from the vision encoder
@@
6:   $\bar{p}^{n},\bar{p}^{a}\leftarrow\sum_{k=1}^{K}\alpha_{k}p_{k}^{n},\sum_{k=1}^{K}\alpha_{k}p_{k}^{a}$triangleright aggregate prompts with learned weights
7:   $T_{n},T_{a}\leftarrow f_{text}(\bar{p}^{n},\bar{p}^{a})$triangleright encode normal/anomaly prompt
@@
11:   $\phi_{c}\leftarrow\operatorname{entmax}_{1.5}\bigl{(}\beta_{c}+\text{noise}\bigr{)}$triangleright ensure sparsity
12:   $T_{att}\leftarrow\sum_{c=1}^{C}\phi_{c}\,u_{c}$triangleright construct attribute representation
@@
14:   $T_{n}^{\rm fus},T_{a}^{fus}\leftarrow\mathrm{Projector}(T_{att},T_{n},T_{a})$triangleright fuse by projector
15:   $\hat{y}=\sigma(sim(x_{\rm cls},T_{(\cdot)}^{fus}))$triangleright image-level score
16:   $\mathcal{L}_{\rm image}=\mathrm{BCE}(\hat{y},y)$triangleright image-level loss
@@
18:    $\hat{G}^{(m)}=\sigma(sim(X^{(m)},T_{(\cdot)}^{\rm fus}))$triangleright pixel anomaly map
@@
20:   $\mathcal{L}_{\rm pixel}=\frac{1}{|M|}\sum_{m\in M}[\mathrm{Focal}(\hat{G}^{(m)},G)+\mathrm{Dice}(\hat{G}^{(m)},G)]$triangleright pixel-level loss
21:   $\mathcal{L}_{\rm PILOT}=\mathcal{L}_{\rm image}+\mathcal{L}_{\rm pixel}$triangleright total loss
@@
4:   $x_{\rm cls}^{t}\leftarrow f_{visual}(I)$triangleright extract global features
5:   $\bar{p}^{n},\bar{p}^{a}\leftarrow\sum_{k=1}^{K}\alpha_{k}p_{k}^{n},\sum_{k=1}^{K}\alpha_{k}p_{k}^{a}$triangleright aggregate prompts with learned weights
6:   $T_{n},T_{a}\leftarrow f_{text}(\bar{p}^{n},\bar{p}^{a})$triangleright encode normal/anomaly prompt
@@
9:   $\phi_{c}\leftarrow\operatorname{entmax}_{1.5}\bigl{(}\beta_{c}+\text{noise}\bigr{)}$triangleright ensure sparsity
10:   $T_{att}\leftarrow\sum_{c=1}^{C}\phi_{c}\,u_{c}$triangleright construct attribute representation
11:   $T_{n}^{\rm fus},T_{a}^{fus}\leftarrow\mathrm{Projector}(T_{att},T_{n},T_{a})$triangleright fuse by projector
12:   $\tilde{y}=\sigma(\mathrm{sim}(x_{\rm cls}^{t},T_{(\cdot)}^{fus}))$triangleright image-level anomaly score
@@
[…]

### v2

@@ 1 Introduction
To address these challenges, we propose a novel framework, Prompt Learning with Integrated Label-Free Test-Time Adaptation for ZerO-ShoT Anomaly Detection (PILOT), a framework designed to mitigate the limitations of current ZSAD methods under domain shift. Instead of relying on a single learnable prompt, PILOT constructs a pool of learnable prompts, allowing the prompts to adaptively learn from various anomalies present in the auxiliary dataset. This design enables the model to adaptively weight the most relevant prompts for each input image, allowing for more nuanced representations of diverse anomalies and supporting more effective TTA. During inference, our label-free test-time adaptation strategy updates all learnable prompts according to their alignment with each test image, using high-confidence pseudo-labels generated from unlabeled target data. This targeted adaptation facilitates rapid and reliable adjustment to new domains (i.e., the test domain) without requiring any labeled data. In addition, to mitigate overfitting to the auxiliary dataset and enhance adaptation stability, PILOT incorporates an attribute memory bank, which consists of a curated set of fixed prompts that describe normal and anomalous states, leveraging pretrained semantic knowledge from the text encoder. This attribute memory bank serves as a semantic anchor, helping to stabilize localization performance during TTA. Together, these innovations enable PILOT to achieve robust anomaly detection and localization across a variety of challenging real-world scenarios. Our contributions are summarized as follows:
@@ 2 Proposed Method
ZSAD involves detecting and localizing anomalies in images from categories not encountered during the training phase. Formally, in the training phase, we employ an auxiliary dataset defined as $\mathcal{D}_{\mathrm{aux}}={(I_{i},y_{i},G_{i})}_{i=1}^{|\mathcal{D}_{\mathrm{aux}}|}$, where each image $I_{i}\!\in\!\mathbb{R}^{H\times W\times 3}$ is paired with an image-level label $y_{i}\!\in\!\{0,1\}$ (0 denotes normal and 1 denotes anomalous) and a pixel-level anomaly mask
$G_{i}\!\in\!\{0,1\}^{H\times W}$. In the inference phase, we consider an unlabeled target dataset $\mathcal{D}_{\mathrm{tar}}=\bigl{\{}I_{t}\bigr{\}}_{t=1}^{|\mathcal{D}_{\mathrm{tar}}|},$ where each $I_{t}\in\mathbb{R}^{H\times W\times 3}$. For every $I_{t}$, the model outputs an anomaly score $\hat{y}_{t}\in[0,1],$ (with values closer to 1 indicating higher abnormality likelihood), along with a $\hat{G}_{t}\in[0,1]^{H\times W}$ that spatially localizes potential anomalous regions at pixel level. Crucially, the category sets of $\mathcal{D}_{\mathrm{aux}}$ and $\mathcal{D}_{\mathrm{tar}}$ are strictly disjoint, ensuring that the target domain remains unseen during training and thus establishing a true zero-shot setting.
@@ 2.1 Learnable Prompt Pool
$p^{n}=[v^{1},\!\dots,\!v^{L},\!\text{``[normal state] object''}],\;p^{a}=[v^{1},\!\dots,\!v^{L},\!\text{``[anomalous state] object''}]$ (1)
@@ 2.3 Prompt Aggregation via Orthogonal Projection
Following [Yang et al.(2021)Yang, He, Fan, Shi, Xue, Li, Ding, and Huang], to robustly leverage both the stable semantics from the attribute embedding $T_{att}$ and the adaptive semantics from the learnable prompts ($T_{n}$ for normal and $T_{a}$ for anomalous), we aggregate these representations using an orthogonal projection operation instead of simple averaging or concatenation. For each prompt embedding $T_{(\cdot)}$ (where $(\cdot)\in\{n,a\}$), we decompose it into components parallel and orthogonal to $T_{att}$:
@@ 3 Experiments
Table 2: Comparison of PILOT vs. state-of-the-art on industrial benchmarks. $\mathbf{Img}$ and $\mathbf{Pix}$ denote image-level and pixel-level metrics, respectively. “w/o TTA” and “w/ TTA” indicate performance without and with TTA, respectively. Arrows indicate change in performance: $\uparrow$(increase >0.5>0.5), $\downarrow$(decrease <0.5<-0.5), and $\circ$(no significant change, $\pm 0.5$).
[…]
@@ 3.1 Comparison with State-of-the-Art
As reported in Table 1, PILOT establishes state-of-the-art performance across all benchmarks for both image-level anomaly detection and pixel-level localization in the zero-shot setting. This can be benefited from PILOT comprehensive approach in addressing critical challenges prevalent in ZSAD, that explicitly tackles domain shift in test phase. Prior approaches [Cao et al.(2023b)Cao, Xu, Sun, Cheng, Du, Gao, and Shen, Jeong et al.(2023)Jeong, Zou, Kim, Zhang, Ravichandran, and Dabeer] rely solely on predefined prompts, which may not generalize well to unseen domains, whereas methods such as [Chen et al.(2023)Chen, Han, and Zhang, Cao et al.(2024)Cao, Zhang, Frittoli, Cheng, Shen, and Boracchi, Zhou et al.(2022a)Zhou, Yang, Loy, and Liu, Zhou et al.(2024)Zhou, Pang, Tian, He, and Chen] employ lightweight visual adapters or a single learnable prompt to capture anomaly cues, but often overfit to the auxiliary training set, limiting robustness under distribution shift. PILOT distinguishes itself by dynamically identifying the most relevant prompts for each input and refining only those learnable prompt parameters during inference using high-confidence pseudo-labels from the unlabeled target data. Note that while existing baselines can utilize the proposed image-level TTA strategies, these often degrade pixel-level localization performance and do not yield consistent improvements across both detection and localization tasks. On the contrary, PILOT uniquely achieves gains in both image-level and pixel-level metrics under TTA, a phenomenon analyzed in detail in Section 3.2. We also compare against state-of-the-art full-shot methods in Appendix F.1, which show that PILOT achieves image-level accuracy on par.
@@ 3.3 Module Ablation
[…]
@@ 4 Conclusion
Acknowledgements
The authors wish to acknowledge the financial support of McGill University, the Natural Sciences and Engineering Research Council of Canada (NSERC), and the Fonds de Recherche du Québec – Nature et technologies (FRQNT). We also acknowledge the computing resources provided by Calcul Québec and the Digital Research Alliance of Canada.
@@ Acknowledgements
This appendix provides supplementary details for the BMVC 2025 paper titled ”Zero-Shot Anomaly Detection with Dual-Branch Prompt Selection”.
@@ B.2 Domain Discrepancy Analysis
To validate that MVTec AD and VisA indeed represent statistically distinct domains suitable for strict zero-shot evaluation, we quantitatively assessed the distributional discrepancy between their CLIP global feature embeddings. Specifically, we employed Maximum Mean Discrepancy (MMD2[Gretton et al.(2012)Gretton, Borgwardt, Rasch, Schölkopf, and Smola], Fréchet Inception Distance (FID) [Heusel et al.(2017)Heusel, Ramsauer, Unterthiner, Nessler, and Hochreiter], and the Kolmogorov–Smirnov (KS) [Massey Jr(1951)] test. Our analysis yielded $\mathrm{MMD}^{2}=0.0328$ ($p$-value =0.0099=0.0099), indicating a significant difference in embedding distributions. The FID score between the two datasets is $51.9$, further corroborating a substantial distributional gap. Additionally, the KS test computed across all feature dimensions revealed that the minimum, mean, and maximum $p$-values are 0, $5.16\times 10^{-5}$, and $0.0329$, respectively, suggesting statistically significant differences in nearly every dimension of the embedding space. These results collectively confirm that MVTec AD and VisA are non-overlapping in the learned representation space, thus justifying their use as auxiliary and target datasets in our zero-shot anomaly detection experiments.
@@ D.1 PILOT Details
Appendix D provides full implementation details of PILOT. We implement PILOT atop the open-source CLIP ViT-L/14@336px backbone, freezing all transformer weights except for the final visual-to-joint projection layer, which is jointly optimized alongside our prompt modules. Our learnable prompt pool $\mathcal{P}$ contains $K=5$ paired prompts that each comprise $L=12$ learnable vectors for “normal” and “anomalous” semantics, while a fixed attribute memory bank (see Appendix J) provides semantic tokens extracted from structured templates. All images are uniformly resized to $518\times 518$ pixels, and we extract both the [CLS] token $x_{\mathrm{cls}}$ and patch embeddings $\{X^{(M)}\}$ at layers 6, 12, 18, and 24 of the CLIP encoder. All prompt embeddings are initialized by sampling from a Gaussian distribution $\mathcal{N}(0,\,0.02)$. PILOT is trained for five epochs using AdamW with a base learning rate of $10^{-3}$, a batch size of 32. During test-time adaptation, we fix the pseudo-label fraction at $\rho=0.25$. All experiments are conducted in PyTorch 2.1.0 on a single NVIDIA Ada A5000 GPU with 32 GB of memory.
@@ E.2 Pseudo label Statistic
In Table 6 we report three statistics that describe how our TTA selection (See Section 2.5 of the main paper) distributes pseudo-labels over the target dataset $D_{\mathrm{tar}}$. Let $B=\lvert D_{\mathrm{tar}}\rvert$ and recall that we form the top-$\rho$ set $A$ and bottom-$\rho$ set $N$, with $S=A\cup N$. For each class $i$ in the set of classes $\mathcal{I}$, we write $g_{i}$ for the fraction of samples in $S$ belonging to class $i$.
@@
E.3 Other TTA Method
@@ E.7 Prompt Aggregation Ablation
[…]
@@ Appendix G Medical anomaly localization results
Medical anomaly localization results are reported in Table 14, with all evaluations conducted without TTA. Since medical datasets lack explicit “normal” examples, our TTA strategy cannot be applied; nevertheless, PILOT still produces competitive localization maps, achieving the highest mean AUROC and the second-best AUPRO across benchmarks. These results underscore the generality of our dual-branch prompt learning framework: even in the absence of adaptation, dynamically weighting prompts with the most relevant anomaly cues deliver state-of-the-art performance in medical domains.
@@ Appendix I Compared with AnovL
Table 16: Comparison between PILOT and AnovL on two benchmarks (MVTec AD and VisA). “w/o TTA” and “w/ TTA” indicate performance without and with test-time adaptation, respectively. Arrows denote change under TTA: $\uparrow$(improvement), $\downarrow$(decline), $\circ$(no significant change, $\pm 0.5$).
[…]
@@
4:   $x_{\rm cls},\{X^{(m)}\}_{m\in M}\leftarrow f_{visual}(I)$ $\triangleright$ extract global & patch features from the vision encoder
@@
6:   $\bar{p}^{n},\bar{p}^{a}\leftarrow\sum_{k=1}^{K}\alpha_{k}p_{k}^{n},\sum_{k=1}^{K}\alpha_{k}p_{k}^{a}$ $\triangleright$ aggregate prompts with learned weights
7:   $T_{n},T_{a}\leftarrow f_{text}(\bar{p}^{n},\bar{p}^{a})$ $\triangleright$ encode normal/anomaly prompt
@@
11:   $\phi_{c}\leftarrow\operatorname{entmax}_{1.5}\bigl{(}\beta_{c}+\text{noise}\bigr{)}$ $\triangleright$ ensure sparsity
12:   $T_{att}\leftarrow\sum_{c=1}^{C}\phi_{c}\,u_{c}$ $\triangleright$ construct attribute representation
@@
14:   $T_{n}^{\rm fus},T_{a}^{fus}\leftarrow\mathrm{Projector}(T_{att},T_{n},T_{a})$ $\triangleright$ fuse by projector
15:   $\hat{y}=\sigma(sim(x_{\rm cls},T_{(\cdot)}^{fus}))$ $\triangleright$ image-level score
16:   $\mathcal{L}_{\rm image}=\mathrm{BCE}(\hat{y},y)$ $\triangleright$ image-level loss
@@
18:    $\hat{G}^{(m)}=\sigma(sim(X^{(m)},T_{(\cdot)}^{\rm fus}))$ $\triangleright$ pixel anomaly map
@@
20:   $\mathcal{L}_{\rm pixel}=\frac{1}{|M|}\sum_{m\in M}[\mathrm{Focal}(\hat{G}^{(m)},G)+\mathrm{Dice}(\hat{G}^{(m)},G)]$ $\triangleright$ pixel-level loss
21:   $\mathcal{L}_{\rm PILOT}=\mathcal{L}_{\rm image}+\mathcal{L}_{\rm pixel}$ $\triangleright$ total loss
@@
4:   $x_{\rm cls}^{t}\leftarrow f_{visual}(I)$ $\triangleright$ extract global features
5:   $\bar{p}^{n},\bar{p}^{a}\leftarrow\sum_{k=1}^{K}\alpha_{k}p_{k}^{n},\sum_{k=1}^{K}\alpha_{k}p_{k}^{a}$ $\triangleright$ aggregate prompts with learned weights
6:   $T_{n},T_{a}\leftarrow f_{text}(\bar{p}^{n},\bar{p}^{a})$ $\triangleright$ encode normal/anomaly prompt
@@
9:   $\phi_{c}\leftarrow\operatorname{entmax}_{1.5}\bigl{(}\beta_{c}+\text{noise}\bigr{)}$ $\triangleright$ ensure sparsity
10:   $T_{att}\leftarrow\sum_{c=1}^{C}\phi_{c}\,u_{c}$ $\triangleright$ construct attribute representation
11:   $T_{n}^{\rm fus},T_{a}^{fus}\leftarrow\mathrm{Projector}(T_{att},T_{n},T_{a})$ $\triangleright$ fuse by projector
12:   $\tilde{y}=\sigma(\mathrm{sim}(x_{\rm cls}^{t},T_{(\cdot)}^{fus}))$ $\triangleright$ image-level anomaly score
@@
17:  $\mathcal{L}_{image}^{TTA}=\frac{1}{|A|+|N|}\sum_{i\in A\cup N}\mathrm{BCE}(\tilde{y}_{i},\text{pseudo-labels})$ $\triangleright$ image-level TTA loss